how to solve 1/2-(1/3-1/7)
Answer by josh_jordan(263)

to solve, start with what's inside the parenthesis, which is 1/3 - 1/7. To subtract two fractions with different denominators, determine what the LCD is. In other words, what is the first number that is divisible by both 3 and 7? 21 is the first number, so, change the denominators of both fractions to 21. Then muliply the numerator and denominator of each fraction by the number of times the denominator of each fraction goes in to 21. 3 goes into 21 7 times, so multiply 1 and 3 by 7. This gives us 7/21. Now, multiply 1/7 by 3, since 7 goes into 21 3 times. This gives us 3/21. Now subtract the two fractions: 7/21 - 3/21 = 4/21. Now, subtract 4/21 from 1/2. Using the same process as earlier, the LCD of both of these fractions is 42. So, change both denominators to 42. Multiply the first fraction by 21, and the second fraction by 2, which gives us: 21/42 - 8/42. Now we can subtract: 21/42 - 8/42 = 13/42